OpenStudy (samara8954):

A line passing through which of the following pairs of coordinates represents a proportional relationship? (1, 3) and (3, 6) (1, 3) and (2, 6) (1, 2) and (2, 3) (1, 1) and (1, 3)

OpenStudy (samara8954):

@Da_Homie

OpenStudy (aripotta):

i guess you can kinda look at this like they're fractions does 1/3 = 3/6? once you simplify 3/6, you get 1/2. and 1/3 does not equal 1/2. so a is not your answer

OpenStudy (aripotta):

does 1/3 = 2/6?

OpenStudy (samara8954):

yes

OpenStudy (aripotta):

there you go :)

OpenStudy (samara8954):

A line passing through which of the following pairs of coordinates represents a proportional relationship? (2, 5) and (4, 6) (3, 6) and (4, 8) (1, 3) and (3, 6) (2, 4) and (5, 6) END STUDY SESSION

OpenStudy (samara8954):

can you help @AriPotta

OpenStudy (coconutjj):

Use the same strategy @AriPotta taught you earlier

OpenStudy (aripotta):

yea, you'd use the same thing as the other question does 2/5 = 4/6?